# ClassContrast: Bridging the Spatial and Contextual Gaps for Node Representations

### v1

@@ 5.1 Node Classification Results
Node2vec (Grover et al. 2016) 0.856$\pm$0.015 0.894$\pm$0.014 0.919$\pm$0.004
GAE (Kipf and Welling 2016) 0.895$\pm$0.165 0.887$\pm$0.084 0.957$\pm$0.012 0.689$\pm$0.384 0.736$\pm$1.090 0.753$\pm$1.297 0.820
VGAE (Kipf and Welling 2016) 0.852$\pm$0.493 0.810$\pm$0.339 0.929$\pm$0.134 0.669$\pm$0.866 0.783$\pm$0.401 0.767$\pm$0.557 0.802
ARVGE (Pan et al. 2018) 0.913$\pm$0.079 0.878$\pm$0.177 0.965$\pm$0.015 0.711$\pm$0.377 0.789$\pm$0.501 0.765$\pm$0.468 0.837
DGI (Veličković et al. 2018b) 0.898$\pm$0.080 0.955$\pm$0.100 0.912$\pm$0.060
G-VAE (Grover et al. 2019) 0.947$\pm$0.011 0.973$\pm$0.006 0.974$\pm$0.004
GNAE (Ahn and Kim 2021) 0.941$\pm$0.063 0.969$\pm$0.022 0.954$\pm$0.019 0.782$\pm$0.829 0.729$\pm$1.083 0.751$\pm$1.067 0.854
VGNAE (Ahn and Kim 2021) 0.892$\pm$0.067 0.955$\pm$0.055 0.897$\pm$0.040 0.703$\pm$0.120 0.733$\pm$0.573 0.789$\pm$0.302 0.828
GIC (Mavromatis et al. 2021) 0.935$\pm$0.060 0.970$\pm$0.050 0.937$\pm$0.030
LINKX (Lim et al. 2021a) 0.934$\pm$0.030 0.935$\pm$0.050 0.801$\pm$0.380 0.758$\pm$0.470 0.857
DisenLink (Zhou et al. 2022) 0.971$\pm$0.040 0.983$\pm$0.030 0.844$\pm$0.190 0.807$\pm$0.400 0.901
DGAE (Fu et al. 2023) 0.958$\pm$0.044 0.972$\pm$0.034 0.978$\pm$0.012 0.757$\pm$0.586 0.681$\pm$1.207 0.683$\pm$1.279 0.838
VDGAE (Fu et al. 2023) 0.959$\pm$0.042 0.978$\pm$0.030 0.970$\pm$0.012 0.850$\pm$0.478 0.761$\pm$0.475 0.813$\pm$0.849 0.889
ClassContrast 0.967$\pm$0.047 0.952$\pm$0.059 0.980$\pm$0.012 0.796$\pm$0.338 0.814$\pm$0.413 0.831$\pm$0.393 0.905
Table 6: Homophily matrices for the SQUIRREL dataset. Every row represents the corresponding homophily ratio of the row’s class. In every row, the highest homophily ratio is marked bold.
ClassContrast Spatial-1 Homophily Matrix
$\mathbf{C}_{1}$ $\mathbf{C}_{2}$ $\mathbf{C}_{3}$ $\mathbf{C}_{4}$ $\mathbf{C}_{5}$
$\mathbf{C}_{1}$ 0.147 0.146 0.154 0.213 0.340
$\mathbf{C}_{2}$ 0.132 0.162 0.173 0.221 0.312
$\mathbf{C}_{3}$ 0.104 0.156 0.194 0.233 0.314
$\mathbf{C}_{4}$ 0.106 0.146 0.176 0.252 0.320
$\mathbf{C}_{5}$ 0.106 0.139 0.178 0.236 0.341
ClassContrast Spatial-2 Homophily Matrix
$\mathbf{C}_{1}$ $\mathbf{C}_{2}$ $\mathbf{C}_{3}$ $\mathbf{C}_{4}$ $\mathbf{C}_{5}$
$\mathbf{C}_{1}$ 0.167 0.170 0.197 0.229 0.236
$\mathbf{C}_{2}$ 0.133 0.192 0.200 0.229 0.247
$\mathbf{C}_{3}$ 0.121 0.168 0.236 0.226 0.250
$\mathbf{C}_{4}$ 0.119 0.158 0.197 0.261 0.266
$\mathbf{C}_{5}$ 0.118 0.152 0.193 0.230 0.307
ClassContrast Domain Homophily Matrix
$\mathbf{C}_{1}$ $\mathbf{C}_{2}$ $\mathbf{C}_{3}$ $\mathbf{C}_{4}$ $\mathbf{C}_{5}$
$\mathbf{C}_{1}$ 0.174 0.187 0.203 0.212 0.223
$\mathbf{C}_{2}$ 0.164 0.194 0.205 0.214 0.224
$\mathbf{C}_{3}$ 0.165 0.188 0.210 0.214 0.223
$\mathbf{C}_{4}$ 0.162 0.187 0.205 0.220 0.226
$\mathbf{C}_{5}$ 0.161 0.186 0.205 0.216 0.231
@@
Features CORA CITESEER PUBMED TEXAS CORNELL WISC. CHAM. SQUIR.
Spatial Only 80.91$\pm$1.64 60.63$\pm$4.35 85.86$\pm$0.32 67.29$\pm$6.29 49.45$\pm$4.23 57.84$\pm$6.14 63.70$\pm$3.05 44.67$\pm$2.55
Context Only 83.13$\pm$1.94 90.44$\pm$1.27 89.70$\pm$0.50 91.35$\pm$4.37 92.71$\pm$4.23 94.71$\pm$2.62 78.81$\pm$1.33 44.08$\pm$1.58
Both (CC) 88.65$\pm$1.25 86.86$\pm$1.91 91.00$\pm$0.45 90.54$\pm$4.80 90.81$\pm$2.90 93.33$\pm$2.95 84.08$\pm$1.55 54.65$\pm$1.27
Ablation Study.
To evaluate the efficacy of our feature vectors, we conducted ablation studies for the node classification task. We employed three submodels: utilizing (1) spatial embeddings, (2) domain embeddings, and (3) both (ClassContrast embeddings). As given in Table 7, we observe that in the CORA, CHAMELEON, SQUIRREL, and PUBMED datasets, using only spatial or domain feature embeddings individually yields satisfactory performance. However, their combination significantly enhances performance in most cases. SQUIRREL ($+10.57$) and CHAMELEON ($+5.27$) experience a significant increase in accuracy in the combined setting. We note that when the ablation study has a large accuracy gap between the spatial and domain only models for a dataset (e.g., in TEXAS), the accuracies of the SOTA models in Table 2 show huge accuracy deviations for the dataset as well (e.g., TEXAS accuracies range from $54.32$ to $90.30$). A possible explanation is that models might individually capture either spatial or contextual information. Consequently, they may be unable to combine these two sets of features to counterbalance the insufficient information present in one of them, leading to diminished accuracy scores. In contrast, the ablation study offers evidence that the ClassContrast approach is resilient to this limitation and, experiences smaller accuracy losses (e.g., $91.35\rightarrow 90.54$ for TEXAS in Table 7).
5.2 Link Prediction Results
In this part, we show the utilization of ClassContrast Embeddings in the link prediction task. We utilize the common setting described in (Zhou et al. 2022) where the datasets were partitioned into training, validation, and test sets with a ratio of $85\%,5\%,$ and $10\%$ respectively.
The model architecture for the link prediction problem consists of three MLP layers. In this framework, for each considered node pairs $(u,v)$, the normalized CC encodings are term-wise multiplied, $h_{u}\cdot h_{v}$, and feed into MLP. We configure the model with a hidden neuron dimension of 16, a learning rate of 0.001, and train it over 100 epochs with a batch size of 128.
Table 5 reports the prediction results. We report the baselines from (Fu et al. 2023) and (Zhou et al. 2022), which use our experiment settings. Out of three homophilic and three heterophilic datasets, ClassContrast outperforms existing models in three datasets, and gets the second best result in one. On the six datasets, ClassContrast reaches the highest mean AUC of 0.905.
Limitations. The effectiveness of ClassContrast heavily depends on the quality and relevance of the landmarks used to define the contextual embeddings. If these are not representative of the actual data distributions, the model’s performance could be adversely affected. Moreover, the computational overhead associated with generating and processing these enhanced embeddings might be prohibitive for extremely large graphs or datasets with high-dimensional features.
@@ 6 Conclusion
In this work, we introduced ClassContrast, a novel method that integrates domain and spatial information from graphs. Our results show that ClassContrast effectively addresses the predictive limitations of using either spatial or domain information alone, yielding significant improvements when both feature sets are informative. Our computationally efficient model has outperformed or achieved highly competitive results compared to state-of-the-art GNNs across multiple graph tasks, including small/large and homophilic/heterophilic settings. For future work, we plan to extend the ClassContrast approach to effectively integrate with GNNs and to apply it to temporal graphs, incorporating both temporal information and node characteristics.
@@
Appendix A Integrating ClassContrast with GNNs
While our primary goal is to develop an approach that synergistically combines spatial and contextual information, we also conducted additional experiments to evaluate how well the ClassContrast features integrate with existing GNN models. In particular, we tested the three classical GNN models, namely GCN, GraphSAGE and GAT, as well as the recent model LINKX. We replaced the node features as initial node embeddings in original GNN models with our ClassContrast features. We implemented a two-layer GNN framework using the Adam optimizer with an initial learning rate of 0.05. The hyperparameters include a dropout rate of 0.5, a weight decay of 5e-4, and 32 hidden channels. The results presented in Figure 3 and Table 4 indicate a noteworthy improvement in GNN performance when utilizing ClassContrast vectors. This enhancement is attributed to the effective embeddings provided by ClassContrast vectors. Furthermore, the observed accuracy gap remains stable. In our future endeavors, we plan to delve deeper into this direction for a more comprehensive exploration.
Figure 4: Performance comparison of three GNN models (GCN, GraphSAGE, GAT) starting with the original node embeddings (blue) and ClassContrast node embeddings (orange) TEXAS dataset.
Appendix B Heterophily and ClassContrast
B.1 Recent Homophily Metrics
Until recently, the prevailing homophily metrics were node homophily (Pei et al. 2019) and edge homophily (Abu-El-Haija et al. 2019; Zhu et al. 2020). Node homophily simply computes, for each node, the proportion of its neighbors that belong to the same class, and averages across all nodes, while edge homophily measures the proportion of edges connecting nodes of the same class compared to all edges in the network. In the past few years, to study heterophily phenomena in graph representation learning, several new homophily metrics were introduced, e.g., class homophily (Lim et al. 2021b), generalized edge homophily (Jin et al. 2022) and aggregation homophily (Luan et al. 2022), adjusted homophily (Platonov et al. 2023) and label informativeness (Platonov et al. 2023). In Table 8, we give these metrics for our datasets. The details of these metrics can be found in (Luan et al. 2023).
@@ B.1 Recent Homophily Metrics
Metric CORA CITES. PUBMED TEXAS CORNELL WISC. CHAM. SQUIR.
$\mathcal{H}_{CC}$-spat-1 0.8129 0.6861 0.7766 0.1079 0.1844 0.2125 0.2549 0.2564
$\mathcal{H}_{CC}$-context 0.1702 0.1949 0.3245 0.2352 0.2409 0.2449 0.2564 0.2057
$\mathcal{H}_{\text{node}}$ 0.8252 0.7175 0.7924 0.3855 0.1498 0.0968 0.2470 0.2156
$\mathcal{H}_{\text{edge}}$ 0.8100 0.7362 0.8024 0.5669 0.4480 0.4106 0.2795 0.2416
$\mathcal{H}_{\text{class}}$ 0.7657 0.6270 0.6641 0.0468 0.0941 0.0013 0.0620 0.0254
$\mathcal{H}_{\text{agg}}$ 0.9904 0.9826 0.9432 0.8032 0.7768 0.6940 0.6100 0.3566
$\mathcal{H}_{\text{GE}}$ 0.1700 0.1900 0.2700 0.3100 0.3400 0.3500 0.0152 0.0157
$\mathcal{H}_{\text{adj}}$ 0.8178 0.7588 0.7431 0.1889 0.0826 0.0258 0.0663 0.0196
LI 0.5904 0.4508 0.4093 0.0169 0.1311 0.1923 0.0480 0.0015
@@
ClassContrast Spatial-1 Homophily Matrix
@@
ClassContrast Context-S Homophily Matrix
@@
ClassContrast Context-I Homophily Matrix
@@
ClassContrast Spatial-1 Homophily Matrix
@@
ClassContrast Spatial-2 Homophily Matrix
@@
ClassContrast Domain Homophily Matrix
@@
$\mathcal{H}_{\text{class}}(\mathcal{G})\!=\!\frac{1}{C\!-\!1}\sum_{k=1}^{C}\bigg{[}h_{k}\!-\!\frac{\big{|}\{v\mid Z_{v,k}\!=\!1\}\big{|}}{N}\bigg{]}_{+}$
where $\text{where }h_{k}\!=\!\frac{\sum_{v\in\mathcal{V},Z_{v,k}=1}\big{|}\{u\mid u\in\mathcal{N}_{v},Z_{u,:}\!=\!Z_{v,:}\}\big{|}}{\sum_{v\in\{v|Z_{v,k}=1\}}d_{v}}$
@@
where $\mathcal{H}_{\text{node}}^{v}$ is the local homophily value for node $v$; $[a]_{+}=\max(0,a)$; $h_{k}$ is the class-wise homophily metric (Lim et al. 2021b); $\mathrm{Mean}_{u}\left(\{\cdot\}\right)$ takes the average over $u$ of a given multiset of values or variables and $S(\hat{A},Z)=\hat{A}Z(\hat{A}Z)^{\top}$ is the post-aggregation node similarity matrix; $D_{c}=\sum_{v:z_{v}=c}d_{v},\bar{p}_{c}=\frac{D_{c}}{2|\mathcal{E}|}$, $p_{c_{1},c_{2}}=\sum_{(u,v)\in\mathcal{E}}\frac{1\{z_{u}=c_{1},z_{v}=c_{2}\}}{2|\mathcal{E}|},c,c_{1},c_{2}\in\{1,\dots,C\}$.
In Table 9 below, we give ClassContrast Homophily matrices which provides detailed insights on the class interactions in CORA dataset. In particular, Spatial-1 matrix interprets as for any class, the nodes likely to form a link with same class node in their one neighborhood. In the following Contextual-Selective and Contextual-Inclusive Homophily matrices, independent of graph distance, we see the positions of attribute vectors similarity to the chosen class landmarks.
B.2 ClassContrast Homophily Matrices
We provide further homophily matrices for one homophilic and one heterophilic graphs. In Table 9, for CORA dataset, we observe in CC Spatial-1, Context-S and Context-I Homophily matrices show the strong homophily behavior in both spatial and contextual aspects. In Spatial-1 matrix the entry $h_{ij}$ represents how likely the nodes in $\mathcal{C}_{i}$ to connect to nodes in $\mathcal{C}_{j}$ among their $1$-neighborhood. The very high numbers in the diagonal shows that most classes likely to connect with their own class, as CORA’s node homophily ratio (0.83) suggest. From the matrix, we have finer information that the class $\mathcal{C}_{3}$ is very highly homophilic. The domain matrices represents the average similarity/closeness to class landmarks. We observe that in both domain matrices, the node feature vectors likely to land close to their own class landmark, and ClassContrast vectors captures this crucial information.
In Table 10, for WISCONSIN dataset, spatial-1 and Spatial-2 matrices show the irregular behavior as the WISCONSIN node homophily ratio (0.09) suggest. The nodes are unlikely to connect their adjacent nodes in the same class. Even in their two neighborhoods, they don’t have many of their fellow classmate for $\mathcal{C}_{4}$ and $\mathcal{C}_{5}$. However, while the other three classes are not connecting their own classmate, they have several common neighbors. Finally, the domain homophily matrix again shows that while they are not very close in the graph, they all share common interests, as every node’s feature vector lands close to their classmates’.

### v2

@@ 4.1 Node Classification Results
[…]
Figure 3: Performance comparison of three GNN models (GCN, GSAGE, GAT) starting with the original node embeddings (blue) and SCNode node embeddings (orange) on the CORA dataset.
Figure 4: Performance comparison of three GNN models (GCN, GraphSAGE, GAT) starting with the original node embeddings (blue) and SCNode node embeddings (orange) on the TEXAS dataset.
4.2 Link Prediction Results
In this part, we show the utilization of SCNode Embeddings in the link prediction task. We utilize the common setting described in (Zhou et al., 2022) where the datasets were partitioned into training, validation, and test sets with a ratio of $85\%,5\%,$ and $10\%$, respectively.
The model architecture for the link prediction problem consists of three MLP layers. In this framework, for each considered node pair $(u,v)$, the normalized SCNode encodings are term-wise multiplied, $h_{u}\cdot h_{v}$, and feed into MLP. We configure the model with a hidden neuron dimension of 16, a learning rate of 0.001, and train it over 100 epochs with a batch size of 128.
Table˜6 reports the prediction results. We report the baselines from (Fu et al., 2023) and (Zhou et al., 2022), which use our experiment settings. Out of three homophilic and three heterophilic datasets, SCNode outperforms existing models in three datasets and gets the second best result in one. On the six datasets, SCNode reaches the highest mean AUC of 0.905.
Ablation Study.
To evaluate the efficacy of our feature vectors, we conducted ablation studies for the node classification task. We employed three submodels: utilizing (1) spatial embeddings, (2) contectual embeddings, and (3) both (SCNode embeddings). As given in Table˜7, we observe that in the CORA, CHAMELEON, and PUBMED datasets, using only spatial or domain feature embeddings individually yields satisfactory performance. However, their combination significantly enhances performance in most cases. CHAMELEON (+5.27+5.27) experiences a significant increase in accuracy in the combined setting. We note that when the ablation study has a large accuracy gap between the spatial and domain-only models for a dataset (e.g., in TEXAS), the accuracies of the SOTA models in Table 2 show huge accuracy deviations for the dataset as well (e.g., TEXAS accuracies range from $54.32$ to $90.30$). A possible explanation is that models might individually capture either spatial or contextual information. Hence, they may be unable to combine these two sets of features to counterbalance the insufficient information present in one of them, leading to diminished accuracy scores. In contrast, the ablation study offers evidence that the SCNode approach is resilient to this limitation and experiences accuracy gain (e.g., $91.35\rightarrow 94.59$ for TEXAS in Table 7). We present further ablation studies on the effect of neighborhood size choice on spatial embeddings, and similarity metric choice on contextual embeddings in Section˜B.4.
@@ 4.2 Link Prediction Results
Features CORA CITESEER PUBMED TEXAS CORNELL WISC. CHAM.
[…]
Effectiveness in smaller training Settings.
Our model leverages both spatial and contextual encodings, where contextual embeddings are derived from class cluster centroids as landmarks, and spatial encodings capture class distributions within k-hop neighborhoods. For these embeddings to be effective, two conditions are particularly important: reliable landmarks and a balanced class distribution. To evaluate the robustness of SCNode under limited supervision, we fix $20\%$ of the nodes as the test set and vary the size of the training set from $5\%$ to $75\%$ across three benchmark datasets: Cora, Citeseer, and PubMed (see Figure˜5). The results show that SCNode achieves stable performance across all datasets once at least $50\%$ of the nodes are included in the training set. It is also worth noting that on PubMed, our model performs noticeably well even under low training ratios. This is because, despite only using $5\%$ of nodes for training, the dataset still provides nearly 1,000 labeled nodes across only three classes. This ensures a sufficient number of representatives per class, enabling SCNode to learn effective contextual embeddings. Consequently, SCNode produces stronger node representations when enough class representatives are available, making the embeddings more effective for supervised learning.
Figure 5: Training set vs Accuracy
Limitations.
A limitation of SCNode lies in its reliance on the quality and relevance of the landmarks used to define the contextual embeddings. If these landmarks fail to adequately represent the underlying data distributions, the model’s performance may suffer. However, this issue could be mitigated by leveraging multiple landmarks or employing contrastive learning in an unsupervised setting to optimize landmark selection. We aim to address this limitation in future work.
@@ 5 Conclusion
In this work, we introduced SCNode, a novel method that integrates spatial and contextual information from graphs. Our results show that SCNode overcomes the predictive limitations of relying solely on either spatial or contextual features, achieving significant performance gains when both are informative. The model is computationally efficient and consistently outperforms or matches SOTA GNNs across diverse graph tasks, including small and large graphs as well as homophilic and heterophilic settings. Furthermore, the plug-and-play design of SCNode vectors highlights their flexibility, providing a powerful enhancement to existing GNN architectures. In future work, we plan to refine SCNode for deeper integration with GNNs and expand its application to temporal graphs by incorporating temporal dynamics alongside node attributes.
Broader Impact Statement
Our SCNode framework advances graph representation learning by addressing the fundamental limitations of GNNs, particularly their struggles with heterophilic graphs, underreaching, and oversquashing. By integrating spatial and contextual information through a novel landmark-based relative positioning approach, SCNode enhances the adaptability of GNNs across diverse datasets. This breakthrough enables more accurate node classification and link prediction, improving applications in bioinformatics, social networks, and financial modeling. Additionally, SCNode embeddings serve as plug-and-play features, significantly boosting existing GNN models’ performance, thereby bridging the gap between homophilic and heterophilic settings. This research contributes to the broader goal of more generalizable and interpretable graph learning models, expanding their impact across scientific and industrial domains.
Acknowledgments
This research was partially supported by the National Science Foundation under grants DMS-2220613 and DMS-2229417. The authors acknowledge the Texas Advanced Computing Center (TACC) at UT Austin for providing computational resources that have contributed to this paper.
@@ Acknowledgments
Appendix A Homophily and SCNode
A.1 Recent Homophily Metrics
Until recently, the prevailing homophily metrics were node homophily Pei et al. (2019) and edge homophily Abu-El-Haija et al. (2019); Zhu et al. (2020). Node homophily simply computes, for each node, the proportion of its neighbors that belong to the same class, and averages across all nodes, while edge homophily measures the proportion of edges connecting nodes of the same class compared to all edges in the network. In the past few years, to study heterophily phenomena in graph representation learning, several new homophily metrics were introduced, e.g., class homophily Lim et al. (2021b), generalized edge homophily Jin et al. (2022) and aggregation homophily Luan et al. (2022), adjusted homophily Platonov et al. (2023a) and label informativeness Platonov et al. (2023a). In Table˜8, we give these metrics for our datasets. The details of these metrics can be found in Luan et al. (2023).
@@ A.1 Recent Homophily Metrics
Metric CORA CITES. PUBMED TEXAS CORNELL WISC. CHAM.
$\mathcal{H}_{SCN}$-spat-1 0.8129 0.6861 0.7766 0.1079 0.1844 0.2125 0.2549
$\mathcal{H}_{SCN}$-context 0.1702 0.1949 0.3245 0.2352 0.2409 0.2449 0.2564
$\mathcal{H}_{\text{node}}$ 0.8252 0.7175 0.7924 0.3855 0.1498 0.0968 0.2470
$\mathcal{H}_{\text{edge}}$ 0.8100 0.7362 0.8024 0.5669 0.4480 0.4106 0.2795
$\mathcal{H}_{\text{class}}$ 0.7657 0.6270 0.6641 0.0468 0.0941 0.0013 0.0620
$\mathcal{H}_{\text{agg}}$ 0.9904 0.9826 0.9432 0.8032 0.7768 0.6940 0.6100
$\mathcal{H}_{\text{GE}}$ 0.1700 0.1900 0.2700 0.3100 0.3400 0.3500 0.0152
$\mathcal{H}_{\text{adj}}$ 0.8178 0.7588 0.7431 0.1889 0.0826 0.0258 0.0663
LI 0.5904 0.4508 0.4093 0.0169 0.1311 0.1923 0.0480
@@
SCNode Spatial-1 Homophily Matrix
@@
SCNode Context-S Homophily Matrix
@@
SCNode Context-I Homophily Matrix
@@
SCNode Spatial-1 Homophily Matrix
@@
SCNode Spatial-2 Homophily Matrix
@@
SCNode Domain Homophily Matrix
@@
$\mathcal{H}_{\text{node}}(\mathcal{G})==\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\mathcal{H}_{\text{node}}^{v}=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\frac{\big|\{u\mid u\in\mathcal{N}_{v},Z_{u,:}=Z_{v,:}\}\big|}{d_{v}}$
@@
$\mathcal{H}_{\text{class}}(\mathcal{G})\!=\!\frac{1}{C\!-\!1}\sum_{k=1}^{C}\bigg[h_{k}\!-\!\frac{\big|\{v\mid Z_{v,k}\!=\!1\}\big|}{N}\bigg]_{+}\text{where }h_{k}\!=\!\frac{\sum_{v\in\mathcal{V},Z_{v,k}=1}\big|\{u\mid u\in\mathcal{N}_{v},Z_{u,:}\!=\!Z_{v,:}\}\big|}{\sum_{v\in\{v|Z_{v,k}=1\}}d_{v}}$
@@
$\mathcal{H}_{\text{agg}}(\mathcal{G})=\frac{1}{\left|\mathcal{V}\right|}\times\left|\left\{v\,\big|\,\mathrm{Mean}_{u}\big(\{S(\hat{A},Z)_{v,u}^{Z_{u,:}=Z_{v,:}}\}\big)\geq\mathrm{Mean}_{u}\big(\{S(\hat{A},Z)_{v,u}^{Z_{u,:}\neq Z_{v,:}}\}\big)\right\}\right|$
@@
$\text{LI}=-\frac{\sum_{c_{1},c_{2}}p_{c_{1},c_{2}}\log\frac{p_{c_{1},c_{2}}}{\bar{p}_{c_{1}}\bar{p}_{c_{2}}}}{\sum_{c}\bar{p}_{c}\log\bar{p}_{c}}=2-\frac{\sum_{c_{1},c_{2}}p_{c_{1},c_{2}}\log p_{c_{1},c_{2}}}{\sum_{c}\bar{p}_{c}\log\bar{p}_{c}}$
where $\mathcal{H}_{\text{node}}^{v}$ is the local homophily value for node $v$; $a]_{+}=\max(0,a)$; $h_{k}$ is the class-wise homophily metric Lim et al. (2021b); $\mathrm{Mean}_{u}\left(\{\cdot\}\right)$ takes the average over $u$ of a given multiset of values or variables and $S(\hat{A},Z)=\hat{A}Z(\hat{A}Z)^{\top}$ is the post-aggregation node similarity matrix; $D_{c}=\sum_{v:z_{v}=c}d_{v},\bar{p}_{c}=\frac{D_{c}}{2|\mathcal{E}|}$, $p_{c_{1},c_{2}}=\sum_{(u,v)\in\mathcal{E}}\frac{1\{z_{u}=c_{1},z_{v}=c_{2}\}}{2|\mathcal{E}|},c,c_{1},c_{2}\in\{1,\dots,C\}$.
In Table˜9 below, we give SCNode Homophily matrices which provides detailed insights on the class interactions in CORA dataset. In particular, Spatial-1 matrix interprets as for any class, the nodes likely to form a link with same class node in their one neighborhood. In the following Contextual-Selective and Contextual-Inclusive Homophily matrices, independent of graph distance, we see the positions of attribute vectors similarity to the chosen class landmarks.
A.2 SCNode Homophily Matrices
We provide further homophily matrices for one homophilic and one heterophilic graphs. In Table˜9, for CORA dataset, we observe in SCN Spatial-1, Context-S and Context-I Homophily matrices show the strong homophily behavior in both spatial and contextual aspects. In Spatial-1 matrix the entry $h_{ij}$ represents how likely the nodes in $\mathcal{C}_{i}$ to connect to nodes in $\mathcal{C}_{j}$ among their $1$-neighborhood. The very high numbers in the diagonal shows that most classes likely to connect with their own class, as CORA’s node homophily ratio (0.83) suggest. From the matrix, we have finer information that the class $\mathcal{C}_{3}$ is very highly homophilic. The domain matrices represents the average similarity/closeness to class landmarks. We observe that in both domain matrices, the node feature vectors likely to land close to their own class landmark, and SCNode vectors captures this crucial information.
In Table˜10, for WISCONSIN dataset, spatial-1 and Spatial-2 matrices show the irregular behavior as the WISCONSIN node homophily ratio (0.09) suggest. The nodes are unlikely to connect their adjacent nodes in the same class. Even in their two neighborhoods, they don’t have many of their fellow classmate for $\mathcal{C}_{4}$ and $\mathcal{C}_{5}$. However, while the other three classes are not connecting their own classmate, they have several common neighbors. Finally, the domain homophily matrix again shows that while they are not very close in the graph, they all share common interests, as every node’s feature vector lands close to their classmates’.